Find the domain of the following:
a)f(x) =| 1 − x^2 | + 1
b) f(x) = √9 − x^2
c) f(x) = 2^(4−x^2)
For a is the domain the set of all real numbers because it is absolute? For b I will get √(x+3) (x-3) but then I am not sure what to do given the sqrt. And for c...I have no idea lol it would be 2^ (x+2) (x-2)? but then I'm not sure.
Thank you for your help :)

Answer by ikleyn(53765)

.
Find the domain of the following:
a)f(x) =| 1 - x^2 | + 1
b) f(x) = sqrt(9 - x^2)
c) f(x) = 2^(4-x^2)

Thank you for your help :)
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

a) f(x) =| 1 - x^2 | + 1      The domain is the set of all real numbers. The entire number line. From -inf to +inf.

b) f(x) = sqrt(9 - x^2)       The domain is the set of real numbers |x| <= 3:   -3 <= x < =3.   The expression under sqrt must be >= 0.

c) f(x) = 2^(4-x^2)           The domain is the set of all real numbers. The entire number line. From -inf to +inf.